### v1

# Annotation Free Semantic Segmentation with Vision Foundation Models

Abstract
Semantic Segmentation is one of the most challenging vision tasks, usually requiring large amounts of training data with expensive pixel-level annotations. With the success of foundation models and especially vision-language models, recent works attempt to achieve zero-shot semantic segmentation while requiring either large scale training or additional image/pixel-level annotations. In this work, we build a lightweight module on top of a self-supervised pretrained vision encoder to align patch features with a pre-trained text encoder. Importantly, we generate free annotations for any semantic segmentation dataset using existing foundation models and train our alignment module cost free. We use CLIP to detect objects and SAM to generate high quality object masks. Our approach can bring language-based semantics to any pre-trained vision encoder with minimal training. Our module is lightweight, uses foundation models as a sole source of supervision and shows impressive generalization capability from little training data with no annotation.
1 Introduction
In the field of computer vision, end-to-end training has long been the de facto standard, where all model parameters are optimized simultaneously for a given task or loss. Recently, large-scale and inexpensive training data has enabled the surge of foundation models [3]. These models have been conversely leveraged in order to avoid the need for both expensive annotations and extensive computational requirements, as previously required by some vision tasks [26, 19, 38, 25, 1]. While obtaining cheap annotations and leveraging pretrained models is straightforward for image classification tasks, it remains a challenge for dense prediction tasks like semantic segmentation, as existing pretrained models lack either in local feature robustness [26, 17] or in semantic awareness [38, 23, 19].
When deploying foundation models for the semantic segmentation task, there are certain properties that are desirable. First, the model has to be promptable with an open set of categories, in order to swiftly deploy it to new datasets or ontologies, without requiring extensive retraining or fine-tuning on pixel-level annotations. In this setting, we assume that there are no available pairs of image samples and pixel-wise labels. Secondly, the foundation models must generalize to different domains, and must be highly discriminative for dense recognition tasks. The open question is how to leverage existing foundation models, composing them to arrive at a unified model that covers all of the desired properties for a specific setting, and thus avoiding costly retraining.
Figure 1: Overview of our method. We propose a novel approach to annotation free semantic segmentation by composition of different foundation models. a) Label Generation: We detect possible object and background categories in an image using a pretrained image-text model (e.g. CLIP, in pink). We select image patches with high similarity to the text representation of the detected categories. Finally, we pass the location of those patches to a mask proposal network (e.g. SAM, in green). b) Visual Features Alignment: We use the generated segmentations and detected categories to align features from any arbitrary, more expressive, image encoder (e.g. DINOv2, in blue) with a pretrained text encoder (with our alignment module, in purple). c) Test-Time Inference: At test time, the newly aligned image encoder projects image features into text space. They are finally classified according to their similarity to the pre-computed vocabulary prototypes.
Foundation models that align images with text descriptions present some of these properties. In particular, Contrastive language-image pretraining (CLIP [26]), trained on a large set of image-text pairs, has shown great success as a baseline for other tasks, such as zero-shot open-set image classification. It is, however, not designed for dense recognition tasks. The text-image alignment process takes the image as a whole and has no explicit notion of object localization or extent. There exists, however, some evidence that CLIP image features can be used to derive a semantic understanding of different regions within an image [43] [25], producing features that are nonetheless sub-optimal for fine dense recognition tasks such as semantic segmentation. On the other hand, recent works show that self-supervised task-agnostic pretraining methods, when applied to a sufficiently large set of unlabelled image data, can produce robust and generic visual features, suitable for any vision task. In particular, DINOv2 [25] features show generalization across different domains and are discriminative for dense recognition tasks. These models are however trained without any explicit notion of semantic meaning.
In this paper, we propose a novel approach to open vocabulary semantic segmentation by composition of different foundations models as building blocks and source of supervision. The result is a model that is both generalizable, and discriminative of semantic meaning and object shape. We propose a light-weight alignment module, which projects the rich features of a self-supervised model to the embedding space of a model trained with text-image pairs. The resulting features are both highly discriminant and are grounded to semantics, while not requiring human-generated segmentation or image annotations. Our method is generic and can be applied to any pretrained image or text encoder.
Specifically, the proposed alignment module is trained in a contrastive manner to align visual features from a DINOv2 encoder with the text embedding space of a CLIP model. In order to supervise the shape and extension of objects, we generate pseudo ground truth masks from readily available images and existing dataset ontologies, by leveraging one more vision foundation model. We query the text-image model with prompts derived from a generic predefined set of labels and use the location of confident activations as prompts to Segment Anything (SAM [19]). We use the resulting masks as the supervision for the alignment module. To account for the noisy pseudo annotations, we deploy a loss function that relies on pretrained text features as anchors while optimizing the visual representations and their alignment with the text features.
At test time, the alignment module projects image patches into text space, and the text-aligned patches are decoded into pixel-wise semantic segmentation features. Rather than relying on a mask proposal network, such as SAM, at test time we compute prototypes for each category of the ontology in text-feature space, and classify each output pixel according to its similarity to each prototype. This allows us to use an arbitrary category ontology to produce dataset-specific semantic segmentations on-the-fly. An overview of our method is shown in Fig. 1. Our alignment module is light and fast to train with only thousands of unannotated images and few training epochs we achieve SOTA results avoiding the need for expensive pretraining.
The main contributions of this paper are: (1) We propose composing pre-trained language-image and self-supervised vision foundation models by means of a lightweight contrastive alignment module. (2) We propose a method to generate semantic segmentation masks as pseudo annotation with zero pixel or image level labels. (3) We apply the composed model to the zero-shot semantic segmentation setting, where we show generalization to never-seen-before semantic segmentation datasets. 4) We achieve state of the art results annotation-free semantic segmentation tasks.
2 Related Work
Vision Foundation Models
Recent advancements in large-scale pre-training have led to powerful generalist models, the so called foundation models [3], transferable to various tasks and domains. In particular, expressive vision-language models emerge when scaling contrastive pre-training on text-image pairs, as shown by CLIP [26, 17]. These models have been since applied to a range of vision problems, including open vocabulary semantic segmentation [10, 34, 34] and image generation [27]. SAM [19] is a powerful image segmentation model, trained iteratively (while providing weak annotations) on one billion images. SAM is not grounded to semantic meaning and, although there was an initial attempt by the authors to align SAM with CLIP text embeddings, no quantitative results were reported. Additionally, SAM provides segments of different granularities, making it difficult to extract object segmentations directly from a single query point.
DINO [6] and DINOv2 [25] are self-supervised image encoders that have shown to have better performance than text-aligned models in dense prediction tasks. Conversely, DINO models are not aligned with text, making their application to Open Vocabulary scenarios not straightforward. GLIP [21], and more recently GLIPv2 [39], aim to unify object detection and vision-language understanding by training on a large number of image-text pairs and bounding boxes. Grounding DINO [22] extends a closed-set detector [38] to the open-set setting by training a vision-language cross-modality decoder for cross-modality fusion. More recently, Zou et al. (2023) [44] proposed a decoder for semantic segmentation, image captioning, and visual question answering, which was pretrained on panoptic segmentation with 104K images and 4M image-text pairs. The decoder uses cross-attention between image and text queries, requiring a close-set ontology to be known in advance, thus making its application to the open-set setting limited. While other foundation models have been proposed, in this work we show an example of foundation models compositionality, where knowledge from different models (CLIP [26], SAM [19] and DINOv2 [25]) are used to provide strong open vocabulary semantic segmentation model.
Open vocabulary semantic segmentation. Semantic segmentation is the task of assigning a semantic label to each pixel in a given image. In zero-shot semantic segmentation a model is expected to provide pixel-level labels for unseen categories [4]. Open vocabulary semantic segmentation can be considered as a generalized zero-shot semantic segmentation problem, one that aims at segmenting an arbitrary set of classes. Zhao et al. (2017) [40] were the first to introduce the open vocabulary setting by learning a joint embedding for pixels features and textual concepts based on WordNet.
CLIP-based open vocabulary semantic segmentation. Recent works have advanced the state-of-the-art in the open vocabulary semantic segmentation problem, by leveraging self-supervised and vision-language models. Some of these approaches still require large-scale training. LSeg[20] matches the pixel embeddings to text embedding of CLIP. OpenSeg [14] learns visual-semantic alignments by aligning each word in a caption to one or few predicted masks. The general trend for open vocabulary methods is however to build on top of CLIP [8, 10, 36, 41, 28] for better generalization with less supervision. A popular line of work [35, 36, 15, 11, 37] rely on training a class-agnostic mask proposal network, leveraging CLIP embeddings for mask classification with additional techniques to strengthen CLIP image patch embeddings. These approaches still need various degrees of pixel-level supervision, which can be expensive to obtain for fine-grained categories. Training on available limited labelled datasets might thus affect their transferability and generalizability.
Open vocabulary semantic segmentation with zero pixel level supervision. One of the first approaches to zero-shot semantic segmentation with no pixel-level supervision is MaskCLIP [43] , which utilizes CLIP to generate annotations and uses them as pseudo-labels for training a segmentation network. Unlike MaskCLIP, we train a lightweight module on top of a frozen pretrained vision encoder, rather than training a complete new network. Other methods modify the CLIP encoder and update it based on a large set of image-text pairs. GroupVit [34] optimizes a hierarchical pixel grouping strategy integrated in a learned ViT model. TCL [7] trains a decoder to ground masks with language based on datasets of 12M and 3M images. Closer to our work, SAM-CLIP [32] finetunes SAM pretrained model by adding an additional CLIP head and then distilling from both CLIP and SAM using unsupervised data of 41M images. This approach tries to bring semantics to SAM but relies on large scale retraining. Our works uses SAM as a segmentation teacher and aligns a more expressive vision encoder with language, hence requiring much less training data and no model retraining. In particular, our method surpasses previous works performance using only 118 thousand unlabeled images.
Training free methods. ReCo [29] builds a dataset of $K$ images for each concept from large-scale unlabelled dataset. Then a nearest neighbour approach is used to produce initial segmentation of a target concept that is then refined by DenseCLIP [28]. CLIP-DIY [33] divides an image into smaller patches that are then classified by CLIP. Patches are then aggregated and transformed to dense prediction using objectness scores from pretrained foreground-background segmentation model. In our work, we rely on SAM to provide accurate masks of detected objects. All training free methods, require a large preprocessing time and usually many steps of refinements, rendering them infeasible. We use the training-free part of our method as a source of pseudo annotations to then produce an efficient model.
Our work is the first that aligns an off the shelf pretrained and frozen vision encoder with another text encoder, at patches level with little training and light-weight alignment module; making our methods readily accessible and usable by practitioners for plug and play annotation free open vocabulary semantic segmentation.
3 Method
General overview.
Semantic segmentation is the task of assigning a semantic meaning to each pixel of an image. Given the recent advancements in large pretrained models, we ask the question of whether these foundation models can be redeployed by composition to the task of semantic segmentation. We consider 3 main models. 1) CLIP [26], pretrained on a million image-text pairs, exhibiting semantic understanding of the image as a whole but not designed for object localization. 2) Purely self-supervised pretrained vision encoders, such as DINOv2 [25], producing features that transfer well to many downstream tasks, and that are consistent across similar objects and parts of objects, but without an explicit link to semantic notions. 3) SAM [19], trained to segment objects or parts of objects, but lacking proper semantic understanding. We design a method that leverages the distinctive properties of those foundation models to enable zero-shot and open vocabulary semantic segmentation.
To achieve this, we propose a two-stage approach: first, we leverage CLIP and SAM to generate pseudo semantic masks for a given vocabulary of classes. In the second stage, we use the predictions of the first stage to train a small alignment module that aligns a frozen off-the-shelf image encoder,  DINOv2 [25] with a pretrained text encoder at the patch level, resulting in a strong self-supervised semantic segmentation model.
3.1 Preliminary
We consider a dataset of images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$ accompanied with a set of possible categories $\{g_{1},g_{2},\dots,g_{L}\}$. A transformer-based image encoder [12] takes as input an image ${im_{i}}\in\mathbb{R}^{C,H,W}$, divided into patches $im_{i}=[\mathbf{im}^{1}_{i},\dots\mathbf{im}^{p}_{i},\dots,\mathbf{im}^{N}_{i}]$ where $N$ is the number of patches. It then extracts a class token and patches embedding $x_{i}=[\mathbf{cls}_{i},\mathbf{x}^{1}_{i},\dots,\mathbf{x}^{p}_{i},\dots,\mathbf{x}^{N}_{i}]$ where $\mathbf{cls}_{i},\mathbf{x}^{p}_{i}\in\mathbb{R}^{D}_{v}$. CLIP [26] provides an image and a text encoder. The text encoder takes as input a text description, e.g. {a photo of a “class”}, and produces a text feature $\mathbf{t}_{i}$ followed by linear projection to $\mathbb{R}^{D}$. Another linear projection layer maps $\mathbf{cls}_{i}$ from the vision encoder to a joint space with the text features $\mathbb{R}^{D}$. CLIP is trained to maximize the cosine similarity between the image class token $\mathbf{cls}_{i}$ and the text representation $\mathbf{t}_{i}$. The patch features $\{\mathbf{x}^{p}_{i}\}$ can be similarly projected onto the same space $\mathbb{R}^{D}$, however we found that the linear projection layer applied directly to the patches features results in features that are sub-optimally aligned with corresponding text features. To overcome this, we follow a procedure similar to Zhou et al. (2022)  [43], where we apply the value projection $V$ of the self attention layer to the input features of the last transformer block, followed by a layer norm, ignoring the key and query. Refer to the Appendix for more details.
3.2 Stage.1: Object detection & masks generation
In this section, we outline our strategy to generate pseudo semantic segmentation labels using  CLIP and SAM. We employ CLIP to recognize categories present in the image. Stage 1.1 extracts query points for SAM mask generation while Stage 1.2 complements this step by labeling SAM’s automatically generated masks. We first describe the detection of categories present in an image. Then, we detail our pseudo mask generation.
High-resolution feature extraction We divide the image it into $C$ crops in a sliding window fashion, process each crop separately with CLIP and rearrange the patches returned from CLIP to construct the features for the full image. This guarantees precise and high quality feature maps for each image. Besides, for each crop $c$, we extract a classification token $\mathbf{cls}_{c}$. We refer to the Appendix for more details.
Defining the set of concerned categories. To detect classes present in an image our method computes the similarity between the classification token for each crop to a set of text features $T$ of a vocabulary. Specifically, all possible labels in a given dataset are considered as the vocabulary, e.g., for COCO-Stuff a set of 171 classes.
Stage 1.1: Querying SAM with CLIP
Object Detection. Each crop is classified with an object category when the object’s text feature $\textbf{t}_{o}$ has the highest similarity to the classification token $\mathbf{cls}_{c}$ among all text features T in the vocabulary. An object is considered as present in the image if it has been assigned to more than a predefined number of crops (see appendix for more details).
Pseudo mask generation. We compute a similarity matrix between the full image feature map and the text features for the detected categories. For each category we select 5 patches with the highest similarity to their corresponding text feature as query points. We feed these query points, along with the original image to SAM and select the mask with the highest confidence (Fig 2).
Stage 1.2: SAM Masks classification
Stage 1.1 may ignore small objects or generate partial masks for an image (Fig 5). Therefore, stage 1.2 performs a complementary pseudo label generation using CLIP and SAM to further boost the performance.
Automatic mask generation. As an alternative to stage 1.1, we rely on all possible masks extracted from the full image using SAM’s automatic mask generation pipeline. We constrain the masks by size and predicted IOU to filter out the low quality and duplicate masks (See appendix).
Mask labelling In order to classify each generated mask by SAM, we compute CLIP’s mean feature embedding corresponding to the area covered by the mask and compute its similarity to the text features of the detected categories in the image. The class with the highest similarity is selected as the pseudo label for the corresponding mask. We refer to the supplementary for comparison of each step results.
Figure 2: Query point selection and mask generation with SAM: Our method selects 5 patches with the highest similarity to the detected object’s text embedding. The coordinates for the center of these patches are forwarded to SAM which generates 3 different segmentation masks for the object. Our method selects the segmentation mask with the highest confidence.
3.3 Stage.2: Lightweight semantic segmentation
The method, as described so far, provides a way to generate fine semantic segmentations from an image dataset with a predefined vocabulary. The generated masks do not cover every object in an image and might be wrong sometimes. It can be also rendered inefficient when constrained with low compute, as it involves querying two foundation models. To overcome this, we are now concerned with how to align any off-the-shelf pretrained vision encoder with text semantics, and with no human supervision.
Alignment Module. We consider the predicted masks along with their corresponding predicted categories, as produced by our method in 3.2, as pseudo labels which we use to train a simple alignment module to map the image patch features to the text embedding. Such a mapping would ground any pretrained vision encoder with language, for dense prediction tasks. Since supervised models may show a bias towards some object categories, we focus on self-supervised pretrained models. Namely, we consider DINOv2 [25], a recent model trained in a fully self-supervised fashion with no text alignment. To overcome the noisy pseudo annotations we rely on 1) frozen pretrained text features as anchors, 2) already discriminative image patches features 3) a loss function that is robust to noise.
Pseudo label generation. Let $\mathbf{D}$ be a dataset of unlabelled images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$. Using the first stage of the pipeline, we extract object masks for each image $im_{i}$ along with their assigned text features $\{m_{i}^{s},t_{s}\}$. The output of the image encoder (e.g. DINOv2) is $\mathbf{x}_{i}=[\mathbf{x}^{1}_{i},\dots,\mathbf{x}^{p}_{i},\dots,\mathbf{x}^{N}_{i}]$ where $N$ is the number of patches, ignoring the $\mathbf{cls}$ token. Using the generated masks and their associated detected categories we assign to each patch $\mathbf{x}^{p}_{i}$ a pseudo label $y^{p}_{i}\in\{1,..,K\}$ where $K$ is the number of detected categories in the dataset $\mathbf{D}$. From the text encoder we extract $K$ text features $T=\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{K}\}$.
Training the alignment module. The pretrained image encoder remains frozen and we optimize a small alignment module $\mathcal{M}$ to map the patch representations $\{\mathbf{x}^{p}_{i}\}$ to the CLIP text embedding space $z^{p}_{i}=\mathcal{M}(\mathbf{x}^{p}_{i})\in\mathbb{R}^{D}$. Note that the method is agnostic to the specific encoder used. As we strive for simplicity, we design our alignment module as one transformer block with multi-head self-attention layer. The self-attention layer allows each patch to attend to other patches in the image. Since we aim for an open-set semantic segmentation, cross-attention with text features [1, 44] is not used as it would require a joint processing of the image features and a closed set of text features at test time.
For notation clarity, we drop the image index and consider only an across-batch patch index $i;i\in\{1,\dots,N*B\}$ where N is the number of patches in an image and $B$ is the batch size. CLIP [16] contrasts the similarity of text features with image class tokens using a cross entropy loss, where a one to one correspondence exists between an image class token and its text features. In our case, we have many image features $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$ extracted from patches of many images and few text features corresponding to detected categories in $\mathbf{D}$. Upon early experiments with CLIP loss, we found it not scalable to image patches and exhibiting poor convergence.
We thus propose to treat each text feature $\mathbf{t}_{k}$ as a prototype of each semantic category. The similarity of a patch feature $z_{i}$ is to be maximized with the corresponding text prototype $\mathbf{t}_{k};y_{i}=k$ and with other patches of the same category, from any image in the same batch. Note that all the feature vectors (text and image patches) are normalized to have unit norm, and the similarities are expressed as a dot product.
Inspired by supervised contrastive loss SupCon [18] that operates on positive and negative pairs, we construct two types of pairs: pairs of image patches (patch-patch pairs: $<\mathbf{z}_{i},\mathbf{z}_{j}>$) and pairs of image patches and text features (patch-text pairs: $<\mathbf{z}_{i},\mathbf{t}_{k}>$). A patch-patch pair $<\mathbf{z}_{i},\mathbf{z}_{j}>$ is considered positive if the patches belong to the same category class $y_{i}=y_{j}$ and negative otherwise. Patch-text pairs $<\mathbf{z}_{i},\mathbf{t}_{k}>$ are positive if $y_{i}=k$. We construct a loss function of two terms operating on the two types of said pairs:
$\ell_{\text{TSupCon}}=\frac{1}{B*N+K}\left(\sum_{k=1}^{K}\ell_{\mathbf{t}}(\mathbf{t}_{k})+\sum_{i}^{B*N}\ell_{\mathbf{im}}(\mathbf{z}_{i})\right),$ (1)
where B is the batch size and N is the number of patches in an image; K is the number of text features. $\ell_{\mathbf{t}}$ is designated for optimizing patch-text pairs
$\ell_{\mathbf{t}}(\mathbf{t}_{k})=\dfrac{1}{N_{k}}\sum_{i:y_{i}=k}\ell_{\mathbf{t}}(\mathbf{z}_{i},\mathbf{t}_{k}),$ (2)
where $N_{k}$ is the number of patches with label $y=k$.
$\ell_{\mathbf{t}}(\mathbf{z}_{i},\mathbf{t}_{y_{i}})=-\mathbf{z}_{i}^{\top}\mathbf{t}_{y_{i}}$ (3)
$+\log\left(\sum_{k=1}^{K}\exp(\mathbf{z}_{i}^{\top}\mathbf{t}_{k})+\sum_{j\neq i}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j})\right).$
The loss $\ell_{\mathbf{t}}(\mathbf{t}_{k})$, defined for each text feature $\mathbf{t}_{k}$, considers all the patches that belong to the category $y=k$ , represented by the text feature $\mathbf{t}_{k}$. The loss is minimized by maximizing the similarity of the concerned patch-text pairs, normalized over all other constructed pairs (patch-patch and patch-text pairs) for a given patch $\mathbf{z}_{i};y_{i}=k$
The loss applied to each patch feature is defined as follows:
$\ell_{\mathbf{im}}(\mathbf{z}_{i})=\frac{1}{N_{y_{i}}}\sum_{l:y_{l}=y_{i}}\left(-\mathbf{z}_{i}^{\top}\mathbf{z}_{l}+\log\sum_{j\neq i}\exp({\mathbf{z}_{i}^{\top}\mathbf{z}_{j})}\right),$ (4)
where $N_{y_{i}}$ is the number of patches in the batch that have the same label as $y_{i}$. This loss term operates on the image patches and it maximizes the similarities of the patches of the same category across different images. This term is a generalization of SupCon [18] to image patches with no data augmentation.
A similar loss function has been proposed for image classification in [2] and it was shown that $\ell_{t}$ (3) formulation can be expressed as a smooth approximation to the maximum function of SupCon and Cross Entropy (CE) loss; in our case of SupCon on patch-text pair and CE on the patch feature with the corresponding text feature as a representative class prototype. This approximation is key to allow a smooth optimization of the different similarities while tolerating possible noisy pairs, as we show in the loss ablation, Sec 4.4.1. We optimize the alignment module $\mathcal{M}$ with $\ell_{\text{TSupCon}}$ (Eq.  1), for a fixed number of epochs.
Deployment. At test time, we extract $\{\mathbf{x}_{i}\}$ features from the image, using the frozen image encoder. The image patches are forwarded through $\mathcal{M}$, $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$, then we compute for each $z_{i}$ the most similar text feature $\mathbf{t}_{k}$, after which we assign to $\mathbf{z}_{i}$ a label $k$. Text features are precomputed using a frozen text encoder. To provide pixel-level predictions we interpolate the similarities to the original image dimension.
4 Experiments
We evaluate our approach on semantic segmentation datasets followed by ablation of our model and loss design.
4.1 Experimental Setup
Implementation Details. We consider Vit-L/14 for both CLIP and DINOv2. We train our alignment module with SGD optimizer, cosine annealing scheduler and a batch size of 5 images. Refer to the Appendix for a detailed description. We use COCO-Stuff [5] images for pretraining our alignment model, composed of only 118k images. We extract pseudo segmentations of COCO-Stuff vocabulary using the first stage of our pipeline, Stage.1 3.2. Then, we train our alignment module on the pseudo annotations for 10 epochs.
Figure 3: Patch-level alignment between image and class. First row shows the original image from Pascal VOC. Second row shows the similarity between the high-resolution patch features from CLIP-L/14 and the image-level label. Third row shows the similarity map after aligning a DINOv2-L/14 model with FMbSeg.
Figure 4: Qualitative results of zero-shot segmentation. We show 2 samples for each one of the considered datasets. The first row shows the ground truth labels. The second row shows the results of FMbSeg.
Figure 5: Qualitative evaluation of Stage.1.1. SAM query points are shown in green stars. Top sub-figure shows instances of correct detected labels and segmentations by Stage.1.1. Below demonstrates its limitations. Small objects, wrongly detected classes (due to ambiguities) and not enough query points to cover all instances. Stage 1.2 alleviates the issue with small objects and incomplete masks since it labels all image masks generated by SAM.
Datasets. Pascal VOC [13] contains 20 foreground classes and everything else is labeled as a background. To isolate the effect of the background class we evaluate on both (PV-20) excluding the background class and PV-21 with the background class. Pascal Context [24] (PC-59) contains 59 classes of objects and stuff. COCO Stuff [5] (CS-171) contains 80 different objects categories and 91 stuff classes. We evaluate on both object categories only (CO- 80) and both objects and stuff categories (CS- 171). We further consider CityScape [9] (City) and ADE20K [42] (ADE). For PV-21, we add different Stuff categories to predict the background class.
Method Training Data mIoU
PV-21 PV-20 PC-59 CO-80 CS-171 City ADE
GroupVit [34] 41M +Labels 50.4 79.7 18.7 27.5 15.3 11.1 9.2
Mask CLIP [43] - 38.8 74.9 23.6 20.6 16.4 12.6 9.8
ReCo [29] - 25.1 57.7 19.9 31.6 14.8 21.1 11.2
TCL [7] 15M + Labels. 55.0 83.2 33.9 31.6 22.4 24.0 17.1
CLIP-S4 [16] 0.12M - - 33.6 - 22.1 - -
SCCLIP [31] - 59.1 - 30.4 - 22.4 - -
CLIP-DIY [33] - 59.0 - 30.4 - - - -
CaR [30] - 67.6 - 30.5 36.6 - -
SAM-CLIP [32] 41M 60.6 - 29.2 31.5 - 17.1
FMbSeg (ours) 0.12M 67.73 85.65 42.72 57.63 29.88 28.36 16.25
Table 1: Semantic Segmentation performance on various datasets. Best method marked in bold and second best in gray, our method is better or on par with SOTA methods.
Compared Methods. We compare with methods that target the alignment of image features with language for open vocabulary semantic segmentation. We divide methods according to the annotation and training they require.
Image Level Annotation: GroupVit [34] and TCL [7]. Annotation free methods: We consider Mask CLIP [43] (the best performing variant), CLIP-S4 [16] and SAM-CLIP [32]. Training free methods: CLIP-DIY [33], SCCLIP [31], CaR [30] and ReCo [29].
Metrics. We consider the widely adopted Mean Intersection over Union (mIoU). We follow the evaluation protocol of TCL [7]. We don’t apply any post-refinement and only one standard template for text features is considered unlike other approaches using 80 different templates.
4.2 Annotation Free Semantic Segmentation Results
Table 1 reports the mIoU on the different datasets. Results of methods in first block are taken from TCL [7] while second block methods are taken from their corresponding papers.
First, training free methods, CLIP-DIY [33], SCCLIP [31], CaR [30] and ReCo [29] perform inferior to our method mostly with a large margin except for CaR [30] on PV-21 which can be attributed to the post-processing of background in CaR. Second, FMbSeg performs best or second best on all datasets even improving over TCL [7] trained on 15M images with image level labels. FMbSeg outperforms TCL by a large margin of 8% on PC-59 and 26% on CO-80. Third, FMbSeg outperforms SAM-CLIP [32] significantly on PV-21 and PC-59 while being inferior on CS-171 and ADE albeit with a small margin. SAM-CLIP [32] trains SAM encoder on 41M images with no code or model available which makes it infeasible for us to evaluate it on the remaining datasets. Nevertheless, we emphasize that our alignment module is light and trained only for a few epochs on almost 400 times less data compared to  [32]. This makes our method a plug and play approach to semantically segment any dataset.
Due to space limits in the main paper, we extensively ablate different ways of leveraging our approach in Appendix. We first consider the stage.1 only 3.2 as a training free method,  FMbSeg Stage.1.1 and  FMbSeg Stage.1.2. We also consider a semi supervised variant where image level labels are provided to the first stage. Stage.1.2 alone results in inferior results to Stage.1.1, however, the results validate that Stage.1.2 helps avoiding some inherit mistakes of Stage.1.1, leading to a boost in performance when combined. Besides, image-level labels can significantly improve the Stage 1 performance leading to SOTA performance without any training. We further ( in Appendix) examine our alignment module (pretrained on COCO-Stuff) when fine-tuned on specific target datasets mentioned in table 1 with no ground-truth annotations and in a self-supervised manner. This fine-tuning does not add a large over-head due to the efficiency of our lightweight alignment module. As shown in Appendix, in this case, the performance systematically exceeds all self-supervised methods on all datasets.
4.3 Qualitative Results
First we inspect the alignment of patch features with vocabulary after we train our alignment module. Fig. 3 shows the patch-level similarity between the image and the image-level label on Pascal VOC. Notice that the newly aligned image encoder produces more consistent similarity scores for the given categories than those produced by CLIP. This shows the efficacy of aligning a better vision encoder with text rather than trying to improve CLIP patches features which typically requires large scale (re)training.
Next, Fig. 4 shows qualitative results of the annotation free segmentation by FMbSeg on the different datasets.
4.4 Ablation
4.4.1 Choice of Loss Function
In the method section 3.3 we propose a loss function, adapted from [2], that applies contrastive learning while treating text features as prototypes. Here we ablate the effectiveness of the two terms: 1) The supervised contrastive loss [18] (SupCon) alone where text features are considered as examples of the corresponding concepts similar to the image patches of a given category. 2) The prototype loss alone Eq (2). Table 2 reports the mIOU on CS-171. We find that SupCon term alone exhibits poor convergence, with long epochs needed, leading to sub-optimal performance. The prototype loss shows a a better performance and faster convergence, probably due to the smooth approximation of Cross Entropy loss and SupCon loss on text-patch pairs. However, when combined with SupCon on pairs of patches, better performance is shown due to more enhancement in the expressiveness of these patches features. Overall, the unique treatment of TSupCon to the text features allows a smoother generalization, better convergence and hence considerably better performance.
4.4.2 Choice of Architecture
We design our alignment module as a single transformer block where multi-head self-attention is applied over image patches. We ablate our choice against other designs, namely a single linear layer and a Multi-layer perceptron (MLP) with GELU activation. Table 2 reports reports the mIOU on CS-171. The differences are not substantial, with MLP achieving better better performance than linear. The transformer block further improves over the MLP.
Module Arch. CS-171
mIoU
Linear 25.32
MLP 26.43
Transformer block 27.61
Alignment Loss CS-171
mIoU
$\ell_{SupCon}$ 24.61
[…]
$\ell_{\text{TSupCon}}$(1) 27.61
Table 2: Left. Comparison of different design choices for our alignment module, a transformer block has a small advantage. Right. Comparison of different losses. SupCon alone is inferior to our full loss TSupCon performing the best.
5 Conclusion and Future Work
Given the impressive performance of vision-language foundation models and their transferability to various downstream tasks, we consider the problem of open vocabulary annotation free semantic segmentation. By composition of different foundation models, namely CLIP and SAM, we extract free pixel-level annotations. We then propose a lightweight alignment module that projects the embedding of any arbitrary pretrained vision encoder to the text encoder space. By pretraining on COCO-Stuff, we obtain a generalist alignment module that can perform smantic segmentation across a wide range of object categories. Alternatively, our method can be deployed as a plug-and-play customized alignment module for any semantic segmentation dataset with zero annotations. We show SOTA results and open a new direction for zero-shot semantic segmentation by foundational models compositionality. As future work, we want to investigate other image encoders and continuous finetuning on new categories. Additionally, we want to explore how our alignment module can improve VLM models that are based on CLIP to further strengthen their object localization capabilities.