### v2

# Fabien Despinoy

Abstract
Semantic Segmentation is one of the most challenging vision tasks, usually requiring large amounts of training data with expensive pixel level annotations. With the success of foundation models and especially vision-language models, recent works attempt to achieve zeroshot semantic segmentation while requiring either large-scale training or additional image/pixel level annotations. In this work, we generate free annotations for any semantic segmentation dataset using existing foundation models. We use CLIP to detect objects and SAM to generate high quality object masks. Next, we build a lightweight module on top of a self-supervised vision encoder, DinoV2, to align the patch features with a pretrained text encoder for zeroshot semantic segmentation. Our approach can bring language-based semantics to any pretrained vision encoder with minimal training. Our module is lightweight, uses foundation models as the sole source of supervision and shows impressive generalization capability from little training data with no annotation.
1 Introduction
Large-scale and inexpensive training data has recently enabled the surge of foundation models in computer vision [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al.]. These models have been employed to avoid expensive annotations and computational requirements of many vision tasks [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al., Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al., Alayrac et al.(2022)Alayrac, Donahue, Luc, Miech, Barr, Hasson, Lenc, Mensch, Millican, Reynolds, et al.]. Leveraging foundation models for semantic segmentation requires the model to produce pixel-wise predictions generalizable to new datasets, domains and ontologies. Therefore, the model must be 1) promptable with an open set of categories and 2) highly discriminative for dense recognition tasks. Consequently, deploying foundation models to obtain cheap annotation for semantic segmentation is challenging as existing models lack either semantic awareness [Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie, Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] or local feature robustness [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Jia et al.(2021)Jia, Yang, Xia, Chen, Parekh, Pham, Le, Sung, Li, and Duerig].
[…]
In particular, there exists some evidence that Contrastive Language-Image Pretraining (CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.]), trained on a large set of image-text pairs, can be used to derive a semantic understanding of different regions within an image [Zhou et al.(2022)Zhou, Loy, and Dai, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.]. However, CLIP features are sub-optimal for dense recognition tasks such as semantic segmentation. On the other hand, self-supervised task-agnostic pretraining methods applied to a sufficiently large set of unlabelled image data have shown great success in producing robust and generic visual features, suitable for any vision task. Specifically, DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] features show generalization across different domains and are discriminative for dense recognition tasks. DINO models are however trained without any explicit notion of semantic meaning  [Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.].
In this paper, we propose a novel approach to open vocabulary semantic segmentation by composition of different foundations models as building blocks and source of supervision. The result is a model that is highly discriminant, generalizable and grounded to semantic meaning and object shape without requiring human-generated segmentation or image annotations. We propose a lightweight alignment module which is trained in a contrastive manner to align visual features from a DINOv2 encoder with the text embedding space of a CLIP model. In order to supervise the shape and extension of objects, we generate pseudo ground truth masks from readily available images and existing dataset ontologies, by leveraging one more vision foundation model, Segment Anything (SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.]).
At test time, the alignment module projects image patches into text space, and the text-aligned patches are decoded into pixel-wise semantic segmentation features. Each pixel is classified according to its similarity to the text embeddings of a target ontology. This enables the model to use any ontology to produce dataset-specific semantic segmentations on-the-fly. An overview of our method coined as FMbSeg is shown in Fig. 1. Our alignment module is light and fast and with only thousands of unannotated images and few training epochs we achieve SOTA results avoiding the need for expensive pretraining.
The main contributions of this paper are: (1) We propose a method to generate semantic segmentation pseudo annotations with zero pixel or image level labels using CLIP and SAM. (2) We propose composing pretrained language-image and self-supervised vision foundation models by means of a lightweight contrastive alignment module. (3) We apply the composed model to the zeroshot semantic segmentation setting, where we show generalization to never-seen-before semantic segmentation datasets. 4) We achieve state of the art results on annotation-free semantic segmentation task.
2 Related Work
Vision Foundation Models
Recent advancements in large-scale pretraining have led to powerful generalist models, transferable to various tasks and domains  [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al.]. In particular, expressive vision-language models emerge when scaling contrastive pretraining on text-image pairs, as shown by ALIGN and CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Jia et al.(2021)Jia, Yang, Xia, Chen, Parekh, Pham, Le, Sung, Li, and Duerig]. These models have been since applied to a range of vision problems, including open vocabulary semantic segmentation [Ding et al.(2022)Ding, Xue, Xia, and Dai, Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] and image generation [Ramesh et al.(2021)Ramesh, Pavlov, Goh, Gray, Voss, Radford, Chen, and Sutskever]. SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] is a powerful image segmentation model, trained iteratively with weak annotations on one billion images. SAM is not grounded to semantic meaning and although there was an initial attempt by the authors to align SAM with CLIP text embeddings, no quantitative results were reported. DINO [Caron et al.(2021)Caron, Touvron, Misra, Jégou, Mairal, Bojanowski, and Joulin] and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] are self-supervised image encoders demonstrating high performance in dense prediction tasks. However, DINO models are not aligned with text, making their application to Open Vocabulary scenarios not straightforward. In this work we show an example of foundation models compositionality, where knowledge from different models (CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.]) are used to provide strong open vocabulary semantic segmentation model.
Open vocabulary semantic segmentation. Zeroshot semantic segmentation requires a model to provide pixel level labels for unseen categories [Bucher et al.(2019)Bucher, Vu, Cord, and Pérez]. Open vocabulary semantic segmentation generalizes this problem aiming at segmenting an arbitrary set of classes. Zhao et al. (2017) [Zhao et al.(2017)Zhao, Puig, Zhou, Fidler, and Torralba] were the first to introduce the open vocabulary setting by learning a joint embedding for pixels and textual concepts based on WordNet.
CLIP-based methods. The general trend for open vocabulary methods is to build on top of CLIP for better generalization with less supervision [Li et al.(2022b)Li, Zhang, Zhang, Yang, Li, Zhong, Wang, Yuan, Zhang, Hwang, et al., Zhang et al.(2022b)Zhang, Zhang, Hu, Chen, Li, Dai, Wang, Yuan, Hwang, and Gao, Chen et al.(2023)Chen, Li, Lim, Torralba, and Zhao, Ding et al.(2022)Ding, Xue, Xia, and Dai, Xu et al.(2023)Xu, Xiong, Ding, and Tu, Zhong et al.(2022)Zhong, Yang, Zhang, Li, Codella, Li, Zhou, Dai, Yuan, Li, et al., Rao et al.(2022)Rao, Zhao, Chen, Tang, Zhu, Huang, Zhou, and Lu]. LSeg[Li et al.(2022a)Li, Weinberger, Belongie, Koltun, and Ranftl] matches the pixel embeddings to text embedding of CLIP. OpenSeg [Ghiasi et al.(2022)Ghiasi, Gu, Cui, and Lin] learns visual-semantic alignments by aligning each word in a caption to one or few predicted masks. A popular line of work rely on training a class-agnostic mask proposal network, leveraging CLIP embeddings for mask classification with additional techniques to strengthen CLIP image patch embeddings [Xu et al.(2022b)Xu, Zhang, Wei, Lin, Cao, Hu, and Bai, Xu et al.(2023)Xu, Xiong, Ding, and Tu, Han et al.(2023)Han, Zhong, Li, Han, and Ma, Ding et al.(2023)Ding, Wang, and Tu, Yu et al.(2023)Yu, He, Deng, Shen, and Chen]. These works still require a various degrees of pixel-level supervision, which can be expensive to obtain for fine-grained categories. Training on limited available labelled datasets might thus affect their generalization.
Training-free methods. ReCo [Shin et al.(2022)Shin, Xie, and Albanie] builds a dataset of $K$ images for each concept from large-scale unlabelled dataset. Then a nearest neighbour approach is used to produce initial segmentation of a target concept that is then refined by DenseCLIP [Rao et al.(2022)Rao, Zhao, Chen, Tang, Zhu, Huang, Zhou, and Lu]. CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni] divides an image into smaller patches that are then classified by CLIP. Patches are then aggregated and transformed to dense prediction using objectness scores from pretrained foreground-background segmentation model. In our work, we rely on SAM to provide accurate masks of detected objects. All training free methods, require a large preprocessing time and usually many steps of refinement, rendering them infeasible. We use the training-free part of our method (section 3.2) as a source of pseudo annotations to produce an efficient model.
Training-based methods without pixel level supervision. Many methods modify the CLIP encoder and update it based on a large set of image-text annotated pairs. GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] optimizes a hierarchical pixel grouping strategy integrated in a learned ViT model. TCL [Cha et al.(2023)Cha, Mun, and Roh] trains a decoder to ground masks with language based on datasets of 12M and 3M images. Closer to our work, MaskCLIP [Zhou et al.(2022)Zhou, Loy, and Dai] utilizes CLIP to generate annotations for training a complete segmentation network from scratch. Unlike this work, we train a lightweight module on top of a frozen pretrained vision encoder. SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] brings semantics to SAM by large-scale fine-tuning and distilation from both SAM and CLIP models with 41M unannotated images. Our work employs SAM as a segmentation teacher and aligns a more expressive vision encoder with CLIP’s pretrained language encoder, hence requiring much less training data and no model retraining. Particularly, our method surpasses previous works performance using only 118 thousand unlabeled images. Our work is the first that aligns off the shelf vision and text encoders, at patches level with little training and a lightweight alignment module; making our methods readily accessible and usable by practitioners for plug and play annotation-free open vocabulary semantic segmentation.
3 Method
General overview.
In this work, we deploy 3 foundation models by composition for the task of semantic segmentation: 1) CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], pretrained on a million image-text pairs, exhibiting semantic understanding of the image as a whole but not designed for object localization. 2) SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.], trained to segment objects or parts of objects, but lacking proper semantic understanding. 3) DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], producing features that transfer well to many downstream tasks, and that are consistent across similar objects and parts of objects, but without an explicit link to semantic notions. We design a method that leverages the distinctive properties of those foundation models to enable zeroshot and open vocabulary semantic segmentation.
To achieve this, we propose a two-stage approach: first, we leverage CLIP and SAM to generate pseudo semantic masks for a given vocabulary of classes. In the second stage, we use the predictions of the first stage to train a small alignment module that aligns a frozen off-the-shelf image encoder,  DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], with a pretrained text encoder at the patch level, resulting in a strong self-supervised semantic segmentation model.
3.1 Preliminary
We consider a dataset of images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$ accompanied with a set of categories in the vocabulary $V=\{class_{1},\dots,class_{L}\}$. CLIP’s transformer-based image encoder [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.] takes as input an image ${\mathbf{im}_{i}}\in\mathbb{R}^{C,H,W}$ divided into patches $\mathbf{im}_{i}=[im^{1}_{i},\dots im^{p}_{i},\dots,im^{N}_{i}]$ and extracts a class token $\mathbf{cls}_{i}$ and patches embedding $\mathbf{x}_{i}=[x^{1}_{i},\dots,x^{p}_{i},\dots,x^{N}_{i}]$. CLIP’s text encoder takes as input a text description {a photo of a “$class_{i}$”} and produces the text feature $\mathbf{t}_{i}$, with $class_{i}$ $\in V$ corresponding to the image label. Visual and textual features represented by $\mathbf{cls}_{i}$ and $\mathbf{t}_{i}$ are separately projected to a joint embedding space $\mathbb{R}^{D}$ and the cosine similarity between them is maximized during CLIP’s training. While the original CLIP architecture discards the patch features $\mathbf{x}_{i}$, they can be projected onto the same space $\mathbb{R}^{D}$. This would enable us to compute the similarity of any category in $V$ with individual patches $\mathbf{x}^{p}_{i}$ and produce a rough localization of the objects in the image. Refer to the Appendix for more details on the specific architecture for projecting patch features $\mathbf{x}_{i}$ to $\mathbb{R}^{D}$.
3.2 Stage 1: Object detection & masks generation
In this section, we outline our strategy to generate pseudo semantic segmentation labels using  CLIP and SAM. We employ CLIP to recognize categories present in the image and SAM for mask generation. We propose two complementary methods for this, Stage 1.1 and Stage 1.2. We provide further details and examples in the Appendix.
Stage 1.1: Querying SAM with CLIP
High-resolution feature extraction We (over-)sample each image into a high resolution, divide it into $C$ crops in a sliding window fashion, process each crop separately with CLIP and rearrange the patches returned from CLIP (section 3.1) for all crops to construct the features for the full image. This guarantees precise and high quality feature maps for each image. Besides, for each crop $c$, we extract a classification token $\mathbf{cls}_{c}$.
Defining the set of concerned categories. To detect classes present in an image our method computes the similarity between the classification token for each crop to a set of text features corresponding to descriptions extracted from vocabulary $V$ (section 3.1). All possible labels in a given dataset are considered as the vocabulary (e.g. a set of 171 classes for COCO-Stuff).
Object presence detection. Each crop $c$ is classified with an object category when the object’s text feature $\textbf{t}_{o}$ has the highest similarity to the classification token $\mathbf{cls}_{c}$ among all descriptions extracted from vocabulary $V$. An object is considered as present in the image if it has been assigned to more than a predefined number of crops (set to 1 in our experiments).
Pseudo mask generation. We compute a similarity matrix between the full image feature map and the text features for the detected categories (Fig. 2). For each category we select 5 patches with the highest similarity as query points. We feed these points along with the original image to SAM and select the mask with the highest confidence.
Stage 1.2: SAM masks classification
Stage 1.1 may ignore small objects or generate partial masks for an image (Fig. 3) due to sub-optimal query points. Stage 1.2 performs a complementary pseudo label generation using CLIP and SAM to further boost the performance. We refer to the supplementary for comparison of the performance for stage 1.1 and stage 1.2.
Automatic mask generation. We retrieve all possible masks extracted from the full image using SAM’s automatic mask generation pipeline. We constrain the masks by size and predicted IOU to filter out the low quality and duplicate masks (See appendix).
Mask labelling In order to classify each generated mask by SAM, we compute CLIP’s mean feature embedding corresponding to the area covered by the mask and compute its similarity to the text features of the detected categories in the image. The class with the highest similarity is selected as the pseudo label for the corresponding mask.
3.3 Stage 2: Lightweight semantic segmentation
Stage 1 extracts segmentation masks from a dataset with a predefined vocabulary. However, the generated masks might be noisy. Additionally, it involves querying two foundation models and might render inefficient when constrained with low compute. To overcome this, we propose to use annotations generated in Stage 1 to align any off-the-shelf pretrained vision encoder with text semantics and with no human supervision.
Alignment module. We consider the generated masks in 3.2 as pseudo labels for training a simple alignment module that maps the image patch features to the text embedding. Such a mapping would ground any pretrained vision encoder with language for dense prediction tasks. Since supervised models may show a bias towards some object categories, we focus on self-supervised pretrained models. Namely, we consider DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], a recent model trained in a fully self-supervised fashion with no text alignment. To overcome the noisy pseudo annotations we rely on 1) frozen pretrained text features as anchors, 2) already discriminative image patches features 3) a loss function that is robust to noise.
[…]
Training the alignment module. The pretrained image encoder remains frozen and we optimize a small alignment module $\mathcal{M}$ to map the patch representations $\{\mathbf{x}^{p}_{i}\}$ to the CLIP text embedding space $z^{p}_{i}=\mathcal{M}(\mathbf{x}^{p}_{i})\in\mathbb{R}^{D}$. Note that the method is agnostic to the specific encoder used. As we strive for simplicity, we design our alignment module as one transformer block with multi-head self-attention layer. The self-attention layer allows each patch to attend to other patches in the image. Since we aim for an open-set semantic segmentation, cross-attention with text features [Alayrac et al.(2022)Alayrac, Donahue, Luc, Miech, Barr, Hasson, Lenc, Mensch, Millican, Reynolds, et al., Zou et al.(2023)Zou, Dou, Yang, Gan, Li, Li, Dai, Behl, Wang, Yuan, et al.] is not used as it would require a joint processing of the image features and a closed set of text features at test time.
For notation clarity, we drop the image index and consider only an across-batch patch index $i;i\in\{1,\dots,N*B\}$ where N is the number of patches in an image and $B$ is the batch size. CLIP [He et al.(2023)He, Jamonnak, Gou, and Ren] contrasts the similarity of text features with image class tokens using a cross entropy loss, where a one to one correspondence exists between an image class token and its text features. In our case, we have many image features $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$ extracted from patches of many images and few text features corresponding to detected categories in $\mathbf{D}$. Upon early experiments with CLIP loss, we found it not scalable to image patches and exhibiting poor convergence.
We thus propose to treat each text feature $\mathbf{t}_{k}$ as a prototype of each semantic category. The similarity of a patch feature $z_{i}$ is to be maximized with the corresponding text prototype $\mathbf{t}_{k};y_{i}=k$ and with other patches of the same category, from any image in the same batch. Note that all the feature vectors (text and image patches) are normalized to have unit norm, and the similarities are expressed as a dot product.
Inspired by supervised contrastive loss SupCon [Khosla et al.(2020)Khosla, Teterwak, Wang, Sarna, Tian, Isola, Maschinot, Liu, and Krishnan] that operates on positive and negative pairs, we construct two types of pairs: pairs of image patches (patch-patch pairs: $<\mathbf{z}_{i},\mathbf{z}_{j}>$) and pairs of image patches and text features (patch-text pairs: $<\mathbf{z}_{i},\mathbf{t}_{k}>$). A patch-patch pair $<\mathbf{z}_{i},\mathbf{z}_{j}>$ is considered positive if the patches belong to the same category class $y_{i}=y_{j}$ and negative otherwise. Patch-text pairs $<\mathbf{z}_{i},\mathbf{t}_{k}>$ are positive if $y_{i}=k$. We construct a loss function of two terms operating on the two types of said pairs:
$\ell_{\text{TSupCon}}=\frac{1}{B*N+K}\left(\sum_{k=1}^{K}\ell_{\mathbf{t}}(\mathbf{t}_{k})+\sum_{i}^{B*N}\ell_{\mathbf{im}}(\mathbf{z}_{i})\right),$ (1)
where B is the batch size and N is the number of patches in an image; K is the number of text features. $\ell_{\mathbf{t}}$ is designated for optimizing patch-text pairs
$\ell_{\mathbf{t}}(\mathbf{t}_{k})=\dfrac{1}{N_{k}}\sum_{i:y_{i}=k}\ell_{\mathbf{t}}(\mathbf{z}_{i},\mathbf{t}_{k}),$ (2)
where $N_{k}$ is the number of patches with label $y=k$.
$\ell_{\mathbf{t}}(\mathbf{z}_{i},\mathbf{t}_{y_{i}})=-\mathbf{z}_{i}^{\top}\mathbf{t}_{y_{i}}$ (3)
$+\log\left(\sum_{k=1}^{K}\exp(\mathbf{z}_{i}^{\top}\mathbf{t}_{k})+\sum_{j\neq i}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j})\right).$
The loss $\ell_{\mathbf{t}}(\mathbf{t}_{k})$, defined for each text feature $\mathbf{t}_{k}$, considers all the patches that belong to the category $y=k$ , represented by the text feature $\mathbf{t}_{k}$. The loss is minimized by maximizing the similarity of the concerned patch-text pairs, normalized over all other constructed pairs (patch-patch and patch-text pairs) for a given patch $\mathbf{z}_{i};y_{i}=k$
The loss applied to each patch feature is defined as follows:
$\ell_{\mathbf{im}}(\mathbf{z}_{i})=\frac{1}{N_{y_{i}}}\sum_{l:y_{l}=y_{i}}\left(-\mathbf{z}_{i}^{\top}\mathbf{z}_{l}+\log\sum_{j\neq i}\exp({\mathbf{z}_{i}^{\top}\mathbf{z}_{j})}\right),$ (4)
where $N_{y_{i}}$ is the number of patches in the batch that have the same label as $y_{i}$. This loss term operates on the image patches and it maximizes the similarities of the patches of the same category across different images. This term is a generalization of SupCon [Khosla et al.(2020)Khosla, Teterwak, Wang, Sarna, Tian, Isola, Maschinot, Liu, and Krishnan] to image patches with no data augmentation.
A similar loss function has been proposed for image classification in [Aljundi et al.(2023)Aljundi, Patel, Sulc, Chumerin, and Reino] and it was shown that $\ell_{t}$ (3) formulation can be expressed as a smooth approximation to the maximum function of SupCon and Cross Entropy (CE) loss; in our case of SupCon on patch-text pair and CE on the patch feature with the corresponding text feature as a representative class prototype. This approximation is key to allow a smooth optimization of the different similarities while tolerating possible noisy pairs. We optimize the alignment module $\mathcal{M}$ with $\ell_{\text{TSupCon}}$ (Eq.  1), for a fixed number of epochs and provide ablation on different loss functions in the Appendix. Deployment. At test time, we extract $\{\mathbf{x}_{i}\}$ features from the image, using the frozen image encoder. The image patches are forwarded through $\mathcal{M}$, $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$, then we compute for each $z_{i}$ the most similar text feature $\mathbf{t}_{k}$, after which we assign to $\mathbf{z}_{i}$ a label $k$. Text features are precomputed using a frozen text encoder. To provide pixel level predictions we interpolate the similarities to the original image dimension.
4 Experiments
4.1 Experimental Setup
Implementation Details. We consider Vit-L/14 for both CLIP and DINOv2. We train our alignment module with SGD optimizer, cosine annealing scheduler and a batch size of 5 images. We use COCO-Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] images for pretraining our alignment model, composed of only 118k images. We extract pseudo segmentations of COCO-Stuff vocabulary using the first stage of our pipeline (section  3.2). Next, we use them to train our alignment module for 10 epochs. Since Stage 1.2 generates many more masks per image compared to Stage 1.1, in order to balance the pseudo annotations of these complementary methods, we randomly sample $75\%$ of annotations from stage 1.2 (see Appendix for ablation).
Figure 2: Patch level alignment between image and class. First row shows images from Pascal VOC. Second row shows the similarity between patch features from CLIP and the text features of the detected category. Third row shows the similarity map after aligning a DINOv2 model with FMbSeg.
Figure 3: Qualitative evaluation of Stage.1.1. SAM query points generated by our method are shown in green stars. Left shows instances of correct segmentations by Stage.1.1. and Right demonstrates its limitations; small objects, wrongly detected classes (due to ambiguities) and not enough query points to cover all instances. Stage 1.2 alleviates the issue with small objects and incomplete masks since it labels the masks generated accurately by SAM.
Figure 4: Qualitative results of zeroshot segmentation. The first row shows the ground truth labels. The second row shows the results of FMbSeg.
Datasets. COCO Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] contains 80 things categories and 91 stuff classes. We evaluate on both things only (CO-80) and things + stuff (CS-171) categories. Pascal VOC [Everingham et al.(2010)Everingham, Van Gool, Williams, Winn, and Zisserman] contains 20 foreground classes and everything else is labeled as background. To isolate the effect of the background class we evaluate on both (PV-20) excluding the background class and (PV-21) with the background class. For PV-21, we add stuff classes from CS-171 as background categories. Pascal Context [Mottaghi et al.(2014)Mottaghi, Chen, Liu, Cho, Lee, Fidler, Urtasun, and Yuille] (PC-59) contains 59 classes of objects and stuff. CityScapes  [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] (City) contains 30 categories from street view images. ADE20K [Zhou et al.(2019)Zhou, Zhao, Puig, Xiao, Fidler, Barriuso, and Torralba] (ADE) contains 150 categories from various indoor and outdoor scenes.
Method Training Data mIoU
PV-21 PV-20 PC-59 CO-80 CS-171 City ADE
GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] 41M +Labels 50.4 79.7 18.7 27.5 15.3 11.1 9.2
Mask CLIP [Zhou et al.(2022)Zhou, Loy, and Dai] - 38.8 74.9 23.6 20.6 16.4 12.6 9.8
ReCo [Shin et al.(2022)Shin, Xie, and Albanie] - 25.1 57.7 19.9 31.6 14.8 21.1 11.2
TCL [Cha et al.(2023)Cha, Mun, and Roh] 15M + Labels. 55.0 83.2 33.9 31.6 22.4 24.0 17.1
CLIP-S4 [He et al.(2023)He, Jamonnak, Gou, and Ren] 0.12M - - 33.6 - 22.1 - -
SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille] - 59.1 - 30.4 - 22.4 - -
CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni] - 59.0 - 30.4 - - - -
CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] - 67.6 - 30.5 36.6 - -
SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] 41M 60.6 - 29.2 31.5 - 17.1
FMbSeg 0.12M 67.73 85.65 42.72 57.63 29.88 28.37 16.25
Table 1: Semantic Segmentation performance on various datasets. Best method marked in bold and second best in gray, our method is better or on par with SOTA methods.
Compared Methods. We compare with methods that target the alignment of image features with language for open vocabulary semantic segmentation. We divide methods according to the annotation and training they require. Image Level Annotation: GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] and TCL [Cha et al.(2023)Cha, Mun, and Roh]. Annotation-free methods: We consider Mask CLIP [Zhou et al.(2022)Zhou, Loy, and Dai] (the best performing variant), CLIP-S4 [He et al.(2023)He, Jamonnak, Gou, and Ren] and SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari]. Training free methods: CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni], SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille], CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] and ReCo [Shin et al.(2022)Shin, Xie, and Albanie].
Metrics. We consider the widely adopted Mean Intersection over Union (mIoU). We follow the evaluation protocol of TCL [Cha et al.(2023)Cha, Mun, and Roh]. We don’t apply any post-refinement and only one standard template {a photo of a “$class$”} for text descriptions is considered unlike other approaches using 80 different templates during evaluation.
4.2 Annotation-Free Semantic Segmentation Results
Table 1 reports the mIoU on the different datasets. Results of methods in first block are taken from TCL [Cha et al.(2023)Cha, Mun, and Roh] while second block methods are taken from their corresponding papers.
First, training free methods, CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni], SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille], CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] and ReCo [Shin et al.(2022)Shin, Xie, and Albanie] perform inferior to our method mostly with a large margin except for CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] on PV-21 which can be attributed to the post-processing of background in CaR. Second, FMbSeg performs best or second best on all datasets even improving over TCL [Cha et al.(2023)Cha, Mun, and Roh] trained on 15M images with image level labels. FMbSeg outperforms TCL by a large margin of 8% on PC-59 and 26% on CO-80. Third, FMbSeg outperforms SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] significantly on PV-21 and PC-59 while being inferior on CS-171 and ADE albeit with a small margin. SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] trains SAM encoder on 41M images with no code or model available which makes it infeasible for us to evaluate it on the remaining datasets. Nevertheless, we emphasize that our alignment module is light and trained only for a few epochs on almost 400 times less data compared to  [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari]. This makes our method a plug and play approach to semantically segment any dataset.
Due to space limits in the main paper, we extensively ablate different ways of leveraging our approach in Appendix. We evaluate stage 1.1 and 1.2 performance as training free methods. We also consider a semi supervised variant where image level labels are provided to stage 1.1 which significantly improves the performance leading to SOTA performance without any training. In Appendix we further examine our alignment module (pretrained on COCO-Stuff) when fine-tuned on specific target datasets mentioned in table 1 with no ground-truth annotations and in a self-supervised manner. This fine-tuning does not add a large over-head due to the efficiency of our lightweight alignment module. As shown in Appendix, in this case, the performance on ADE20k increases to 21.57%.
4.3 Qualitative Results
First we inspect the alignment of patch features with vocabulary after we train our alignment module. Fig. 2 shows the patch level similarity between the image features and the image level text features on Pascal VOC. Notice that the newly aligned image encoder produces distinctly more consistent similarity heatmaps for the given categories than those produced by CLIP. This shows the efficacy of aligning a better vision encoder with text rather than trying to improve CLIP patches features which typically requires large scale (re)training.
Next, Fig. 4 shows qualitative results of the annotation-free zeroshot segmentation by FMbSeg on the different datasets.
5 Conclusion and Future Work
Given the impressive performance of vision-language foundation models and their transferability to various downstream tasks, we consider the problem of open vocabulary annotation-free semantic segmentation. By composition of different foundation models, namely CLIP and SAM, we extract free pixel level annotations. We then propose a lightweight alignment module that projects the embedding of any arbitrary pretrained vision encoder to the text encoder space. By pretraining on COCO-Stuff, we obtain a generalist alignment module that can perform semantic segmentation across a wide range of object categories. Alternatively, our method can be deployed as a plug-and-play customized alignment module for any semantic segmentation dataset with zero annotations. We show SOTA results and open a new direction for zeroshot semantic segmentation by foundational models compositionality. As future work, we want to investigate other image encoders and continuous fine-tuning on new categories. Additionally, we want to explore how our alignment module can improve VLM models that are based on CLIP to further strengthen their object localization capabilities.